OpenStudy (heart_offire):

Write an equation of the line that passes through (1,5) and (-1, 3). A) y = x + 9 B) y = x + 4 C) y = -x + 4 D) y = -x + 9

OpenStudy (trojanpoem):

slope = dy/dx = 3 - 5/ -1 - 1 = -2/-2 = 1 y = mx + b ( m = 1) y = x + b 3 = -1 + b b = 4 y = x + 4

OpenStudy (heart_offire):

thanks